# Direct Superpoints Matching for Fast and Robust Point Cloud Registration

### v1

@@ 3.3 Loss Functions
Transformation Loss. We apply a L1 loss on the predicted transformed locations of all keypoints with the predicted and ground truth transformation matrix.
@@
Overlap Loss. Inspired by [64], we estimate the overlap values $\mathbf{\hat{O}_{\mathbf{X^{\prime}}}}$ and $\mathbf{\hat{O}_{\mathbf{Y^{\prime}}}}$ using a separate MLP layer based on the enhanced feature $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ and $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$, respectively. The overlap estimation is formulated as a binary classification problem, so we use binary cross-entropy loss:
$\mathcal{L}^{X}_{o}=-\frac{1}{M^{\prime}}\sum_{i}^{M^{\prime}}{o_{x,i}^{*}\cdot\log\hat{o}_{x,i}^{*}+(1-o_{x,i}^{*})\cdot\log\left(1-\hat{o}_{x,i}\right)},$ (4)
where $\hat{o}_{x,i}$ is the estimated overlap probability and $o_{x,i}^{*}$ is the ground truth label. We compute the overlap loss $\mathcal{L}^{Y}_{o}$ for the target point cloud similarly.
@@
The final loss is a weighted sum of all the losses above with
@@
4 Experiments
We evaluate our approach on two benchmarks with overlap ranging from 10% to 75%. The first dataset we evaluate is on ModelNet with two benchmarks settings following [24, 64] and the second dataset is 3DMatch [68] on two benchmarks following [24, 64, 44, 66].
4.1 Implementation details
Our approach is implemented using the PyTorch framework [41] on a system with an Intel i9-1300K CPU and a single RTX 3090 GPU. The network training is performed with the AdamW optimizer [36], utilizing a learning rate of 0.0001 and a weight decay of 0.0001. For the ModelNet dataset, the network is trained for 400 epochs with a batch size of 4. On the 3DMatch dataset, the network is trained for 50 epochs with a batch size of 4 as well. Training the network requires approximately 22 hours on ModelNet and around 2 days on 3DMatch.
4.2 ModelNet and ModelLoNet Benchmarks
The ModelNet40 [58] dataset comprises of synthetic CAD models. Following the data setting in [24, 64], the point clouds are randomly sampled from mesh faces of the CAD models, cropped and subsampled. For the ModelNet and ModelLoNet benchmarks, the average overlap is 73.5% and 53.6%, respectively. Our network is trained exclusively on ModelNet, and we evaluate its generalization performance on ModelLoNet. For benchmarking the performance of our model we use the Relative Rotation Error (RRE) and Relative translation Error (RTE) as the primary metrics. Following [64], we also calculate the Chamfer Distance (CD) between the registered scan pairs.
Table 1: Registration results on the ModelNet and ModelLoNet benchmarks.
@@ 4.2 ModelNet and ModelLoNet Benchmarks
PointNetLK 29.725 0.297 0.0235 48.567 0.507 0.0367
OMNet 2.947 0.032 0.0015 6.517 0.129 0.0074
DCP-v2 11.975 0.171 0.0117 16.501 0.300 0.0268
RPM-Net 1.712 0.018 0.00085 7.342 0.124 0.0050
Predator 1.739 0.019 0.00089 5.235 0.132 0.0083
RegTR 1.473 0.014 0.00078 3.930 0.087 0.0037
Geotransformer 1.568 0.018 - 3.809 0.102 -
@@
The results are shown in Table 1. We compare against correspondence-based approaches [24, 64], coarse-to-fine registration approaches [44], and end-to-end methods [2, 63, 59]. We see that our approach performs well on both benchmarks. Moreover, the low chamfer error suggests that predicted correspondences have very high accuracy. Our approach is also able to outperform methods using post-processing steps like RANSAC [24] by a significant margin, which further strengthens the point that direct superpoints matchings can actually work well for point cloud registration.
4.3 3DMatch and 3DLoMatch Benchmarks
@@
We compare our approach against several approaches including learned corresponding based algorithms [21, 13, 4, 24] and coarse-to-fine approaches [66, 44]. Since the algorithms which use some post-preprocessing steps like RANSAC perform well only with a large number of interest points, we only show results for the maximum number of sampled points (5000). We also compare against several approaches designed to avoid RANSAC [10, 59, 12].
Table 2: Registration results on the 3DMatch and 3DLoMatch benchmarks.
@@
The quantitative results are shown in Table 2. For the 3DMatch benchmark, our approach outperforms all the previous methods. This implies that in cases of significant overlap (>30%), the superpoint correspondences are very distinctive and accurate. One of the problems associated with superpoint matching in point clouds is the resolution issue, where the correspondences might not actually coincide with each other when transformed with the ground truth transformation matrix due to the subsampling of the point cloud. In our case, the resolution issue is automatically handled by the correlation weights. We also show systematic comparisons with other approaches in Table 3. We can see our model runs fast and is compact.
Table 3: Comparisons of inference speed and number of parameters on the 3DMatch dataset.
Methods Matching e2e Time Param
CoFiNet coarse-to-fine 1.922s 5.5M
GeoTransformer coarse-to-fine 0.088s 9.8M
RegTR 0.063s 11.8M
Ours direct 0.073s 7.8M
For the 3DLoMatch benchmark, in comparison to all the approaches that do not use any post-processing, our method performs significantly better than the other approaches. Note that [44] uses LGR as the refinement step to get the best results, but when using all the predicted correspondences with their scores and applying weighted SVD, they get much worse results. This implies that the final set of all the correspondences is not optimal and contains outliers. In our case, these outliers are filtered by the overlap and correlation values, thus providing accurate correspondences. The validity of the correspondences predicted by our approach can be verified by comparing the mean Inlier Ratio in Table 2. We get the highest mean IR, almost 19% better than the second-best approach on 3DMatch and 14% higher in 3DLoMatch.
4.4 Ablation Studies
Table 4: Comparisons of different outlier filtering methods.
Outlier Filtering Method 3DMatch 3DLoMatch
RRE RTE RR RRE RTE RR Time
No filtering 1.462 0.046 93.3 2.652 0.074 64.6 0.073
Ours + RANSAC-50k 2.788 0.090 82.3 4.653 0.115 32.0 0.141
Ours (top 15%) + RANSAC-50k 1.701 0.051 93.0 2.696 0.078 64.5 0.153
Iterative refinement 1.404 0.043 93.8 2.592 0.071 65.7 0.078
@@ 4.4 Ablation Studies
Refer to caption Refer to caption
Figure 3: Left: Effect of increasing the number of attention layers. Right: Comparisons of outlier filtering strategies.
Effectiveness of the feature enhancement. The feature enhancement module is essential in our approach. To assess its effectiveness, we trained various networks with attention layers ranging from 0 to 6. Figure 3 illustrates the performance changes as we increase the number of attention layers. When no feature enhancement is used, the model’s performance is significantly poorer, which is understandable since the source point cloud’s superpoint features lack information about the target point cloud. Performance stabilizes at around 6 attention layers. It is worth noting that increasing the number of layers also leads to higher computation burden.
Using correlation scores for filtering out outliers. To filter out outliers in the predicted correspondences we primarily use the correlation values along with overlap values as weights for weighted SVD. Since these weights are much lower for outlier values, the registration results are very accurate. If we use unweighted SVD (with equal weights for all the correspondences), the performance drops rapidly with the increase in number of correspondences, as shown in Fig. 1.
We experiment with a simple iterative refinement scheme proposed in [44], where we can iteratively re-estimate the transformation by pruning out outliers. We found that the performance of this refinement saturated after 5 iterations. See Table 4 row 4 for results. Fig. 1 also shows that this approach is quite stable in pruning outliers on different number of correspondences and improves the results by a slight margin.
Table 5: Effectiveness of different loss terms.
@@
We also experiment with RANSAC on the set of superpoints correspondences, See Table 4 row 2, 3. But as expected, RANSAC gives slightly worse results while taking more computation time. This implies that using our approach is much more effective in pruning outliers thus eliminating the need for refinement.
@@
Effectiveness of the loss terms. Lastly, we analyze the effectiveness of each loss function. Table 5 shows the results with different loss function configurations. We see that by just using Feature Loss with Transformation loss, the model is able to achieves good performance and using Overlap loss is the final step that helps the model prune remaining outliers. Row 3 in Table 5 again implies that superpoint correspondences are very distinctive.
Figure 4: Qualitative Results on 3Dmatch dataset: The source cloud is shown in Green with the red color representing the matched superpoints. Similarily target cloud is in red color with matching superpoints in green color.
5 Limitations and Broader Impacts
We have identified two limitations in the current architecture, First, our approach requires a decent amount of superpoints in the low overlapping region to perform well. Second, our approach samples about 200 best correspondences, but for the low overlapping region only a few of them lie on the overlapping region thus making it hard for the network to achieve good performance on 3DLoMatch benchmark. Instead of regularly downsampling, structure-aware sampling strategies are needed.
This research has the potential to significantly impact the fields of 3D computer vision and robotics, as well as any domains that rely on accurate and efficient point cloud registration, such as autonomous navigation, 3D mapping, and augmented reality. By introducing a simple yet effective approach for directly matching superpoints in an end-to-end manner, the research not only speeds up the registration process but also enhances its robustness by effectively filtering out outliers without any postprocessing refinement.
6 Conclusion
We introduce a straightforward yet highly effective method for point cloud registration that offers simplicity in understanding and implementation. Our approach directly matches superpoints to establish correspondences, enabling the computation of rigid transformations and correlation weights. To enhance performance in low overlap regions, we augment the correlation weights with predicted overlap values. Experimental evaluations on the 3DMatch and ModelNet datasets demonstrate the efficacy of our approach across different overlap scenarios. Notably, our approach is computationally efficient, robust to outliers, and does not require post-processing steps, making it suitable for real-time applications. In the future, we plan to extend its application to cross-modalities such as image-to-image and image-to-point cloud registration.
Appendix A Appendix
A.1 Network Architecture
Figure 5: Network Architecture of the KPConv backbone for 3DMatch and ModelNet dataset. We do not downsample the point cloud by a large factor to maintain resolution on the overlapping region of the point cloud pair.
Figure 6: Network architecture of the Feature Enhancement Module. The architecture is shown for just one block of self-attention and cross-attention layers but can be extended similarly for improvement in performance.
The network architecture for the KPConv backbone used is shown in Fig. 5 and the architecture for the Feature Enhancement module is shown in Fig. 6.
We do not downsample the point cloud by a large factor in the bakbone unlike other methods [44, 64] as this leads to a very low resolution in the overlapping region of the point cloud pair. Since, we are training the network to learn correlations between the downsampled superpoints, it is beneficial to have a moderately larger number of points.
A.2 Evaluation Metrics
We report Registration Recall (RR), Relative Rotation Error (RRE), Relative Translation Error (RTE), Feature Matching Recall (FMR), and Inlier Ratio (IR) for the 3DMatch dataset following common practice [44, 64, 24, 66, 4]. For ModelNet dataset, we report RRE, RTE, and Chamfer Distance Error (CD) following [64, 24, 59].
A.2.1 3DMatch/3DLoMatch
Inlier Ratio is the fraction of inlier correspondences in all predicted correspondences. Corresponding points are considered inliers if the distance between them under the ground-truth transformation is smaller than $\tau_{1}=10cm$
@@ A.2 Evaluation Metrics
Feature Matching Recall is the fraction of point cloud pairs whose inlier ratio is above $\tau_{2}=0.05$. It indicates the likelihood of recovering the optimal transformation between two point clouds. It is a good indicator for methods which use post-processing refinement steps like RANSAC [66, 44, 24] as having a higher FMR increases the probability of recovering the correct transformation. If the total number of point cloud pairs in the test set is $\mathcal{N}$, FMR can be calculated as
$\textnormal{FMR}=\frac{1}{\mathcal{N}}\sum^{\mathcal{N}}_{i=1}\llbracket\textnormal{IR}>\tau_{2}\rrbracket.$ (8)
Registration Recall represents the fraction of point cloud pairs whose Root Mean Square Error (RMSE) is less than $\tau_{3}=20cm$. RR directly measures the effectiveness of the registration algorithm.
$\textnormal{RR}=\frac{1}{\mathcal{N}}\sum^{\mathcal{N}}_{i=1}\left\llbracket\sqrt{\frac{1}{|\mathcal{C^{*}}|}\sum_{(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathcal{C^{*}}}\lVert\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(\mathbf{x}_{i})-\mathbf{y}_{i}\rVert^{2}_{2}}<\tau_{3}\right\rrbracket,$ (9)
where $\mathcal{C^{*}}$ is the set of ground truth correspondences and $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}$ is the predicted transformation matrix.
Relative Rotation Error and Relative Translation Error are the rotation and translation errors between the ground truth transformation matrix $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{T}\mkern-1.5mu}\mkern 1.5mu^{\mathbf{X}}_{\mathbf{Y}}\in SE(3)$ and the predicted transformation $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}\in SE(3)$. We can obtain $\mathbf{R}\in SO(3)$, $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{R}\mkern-1.5mu}\mkern 1.5mu\in SO(3)$, $\mathbf{t}\in\mathbf{R}^{3}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{t}\mkern-1.5mu}\mkern 1.5mu\in\mathbf{R}^{3}$ as the predicted rotation matrix, ground truth rotation matrix, predicted translation vector and ground truth translation vector respectively. Thus, RRE and RTE can be calculated as follows
$\textnormal{RRE}=\textnormal{arccos}\left(\frac{\textnormal{trace}(\mathbf{R}^{T}\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{R}\mkern-1.5mu}\mkern 1.5mu)-1}{2}\right),~{}~{}~{}\textnormal{RTE}=\lVert\mathbf{t}-\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{t}\mkern-1.5mu}\mkern 1.5mu\rVert.$ (10)
A.2.2 ModelNet/ModelLoNet
Chamfer Distance measures the dissimilarity or discrepancy between two sets of points or point clouds. It quantifies the minimum average distance between each point in one set and its nearest neighbor in the other set. It provides a measure of how well two point clouds align or match with each other, with lower values indicating a better alignment.
$\text{{CD}}(\mathbf{X},\mathbf{Y})=\sum_{x\in\mathbf{X}}\min_{y\in\mathbf{Y}}\|y-\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(x)\|_{2}^{2}+\sum_{y\in\mathbf{Y}}\min_{x\in\mathbf{X}}\|y-\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(x)\|_{2}^{2},$ (11)
where $\mathbf{X}$ and $\mathbf{Y}$ represent the sets of points in two point clouds, $\|\cdot\|_{2}$ denotes the Euclidean distance and $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}$ is the predicted transformation between point clouds $\mathbf{X}$, and $\mathbf{Y}$.
A.3 Additional Results
Table 6 shows the scene-wise registration results on the 3DMatch and 3DLoMatch benchmarks. Following literature [44, 64, 24], we only calculate mean RRE and RTE for the successfully registered scan pairs. For the 3DMatch benchmark, our method outperforms other methods in every metric. Our methods achieves not only the highest registration recall, but also the lowest mean RRE and RTE. This implies that the for all the successfully registered points clouds, the predicted correspondences and their correlation values are very accurate. For the 3DLoMatch benchmark, our method outperforms every method which does not use post-processing refinement like RegTR, while giving the lowest RRE and RTE values. In comparison to approaches which use post-processing refinement like GeoTransformer, Predator, and CoFiNet, the performance gap is expected as the predictions in 3DLoMatch contain a much higher number of outliers. But the gap is minor and our approach runs faster.
Table 6: Scene-wise registration results on the 3DMatch and 3DLoMatch benchmarks.
Method 3DMatch 3DLoMatch
Kitchen Home 1 Home 2 Hotel 1 Hotel 2 Hotel 3 Study Lab Mean Kitchen Home 1 Home 2 Hotel 1 Hotel 2 Hotel 3 Study Lab Mean
Registration Recall (%)normal-$\uparrow$
3DSN 90.6 90.6 65.4 89.6 82.1 80.8 68.4 60.0 78.4 51.4 25.9 44.1 41.1 30.7 36.6 14.0 20.3 33.0
FCGF 98.0 94.3 68.6 96.7 91.0 84.6 76.1 71.1 85.1 60.8 42.2 53.6 53.1 38.0 26.8 16.1 30.4 40.1
D3Feat 96.0 86.8 67.3 90.7 88.5 80.8 78.2 64.4 81.6 49.7 37.2 47.3 47.8 36.5 31.7 15.7 31.9 37.2
Predator 97.6 97.2 74.8 98.9 96.2 88.5 85.9 73.3 89.0 71.5 58.2 60.8 77.5 64.2 61.0 45.8 39.1 59.8
CoFiNet 96.4 99.1 73.6 95.6 91.0 84.6 89.7 84.4 89.3 76.7 66.7 64.0 81.3 65.0 63.4 53.4 69.6 67.5
RegTR 97.8 90.6 75.5 97.8 94.9 100 88.5 91.1 92.0 66.2 58.5 64.9 72.7 61.3 70.7 53.0 71.0 64.8
GeoT 98.9 97.2 81.1 98.9 89.7 88.5 88.9 88.9 91.5 85.9 73.5 72.5 89.5 73.2 66.7 55.3 75.7 74.0
Ours 97.6 88.7 87.4 97.8 93.6 100 87.2 97.8 93.7 62.3 55.3 69.4 72.2 62.0 45.6 46.2 76.8 65.0
Relative Rotation Error (°)normal-$\downarrow$
3DSN 1.926 1.843 2.324 2.041 1.952 2.908 2.296 2.301 2.199 3.020 3.898 3.427 3.196 3.217 3.328 4.325 3.814 3.528
FCGF 1.767 1.849 2.210 1.867 1.667 2.417 2.024 1.792 1.949 2.904 3.229 3.277 2.768 2.801 2.822 3.372 4.006 3.147
D3Feat 2.016 2.029 2.425 1.990 1.967 2.400 2.346 2.115 2.161 3.226 3.492 3.373 3.330 3.165 2.972 3.708 3.619 3.361
Predator 1.861 1.806 2.473 2.045 1.600 2.458 2.067 1.926 2.029 3.079 2.637 3.220 2.694 2.907 3.390 3.046 3.412 3.048
CoFiNet 1.910 1.835 2.316 1.767 1.753 1.639 2.527 2.345 2.011 3.213 3.119 3.711 2.842 2.897 3.194 4.126 3.138 3.280
RegTR 1.729 1.34 1.79 1.639 1.28 1.810 1.57 1.35 1.56 3.366 2.446 3.244 2.732 2.439 2.919 3.044 2.428 2.827
GeoT 1.79 1.353 1.79 1.52 1.328 1.57 1.952 1.678 1.625 2.35 2.30 2.541 2.455 2.49 2.504 3.01 2.716 2.547
Ours 1.780 1.272 1.596 1.457 1.239 1.456 1.492 1.196 1.436 2.330 2.281 2.59 2.72 2.520 2.58 2.775 2.61 2.55
Relative Translation Error (m)normal-$\downarrow$
3DSN 0.059 0.070 0.079 0.065 0.074 0.062 0.093 0.065 0.071 0.082 0.098 0.096 0.101 0.080 0.089 0.158 0.120 0.103
FCGF 0.053 0.056 0.071 0.062 0.061 0.055 0.082 0.090 0.066 0.084 0.097 0.076 0.101 0.084 0.077 0.144 0.140 0.100
D3Feat 0.055 0.065 0.080 0.064 0.078 0.049 0.083 0.064 0.067 0.088 0.101 0.086 0.099 0.092 0.075 0.146 0.135 0.103
Predator 0.048 0.055 0.070 0.073 0.060 0.065 0.080 0.063 0.064 0.081 0.080 0.084 0.099 0.096 0.077 0.101 0.130 0.093
CoFiNet 0.047 0.059 0.063 0.063 0.058 0.044 0.087 0.075 0.062 0.080 0.078 0.078 0.099 0.086 0.077 0.131 0.123 0.094
RegTR 0.040 0.041 0.057 0.057 0.042 0.039 0.054 0.058 0.049 0.079 0.064 0.078 0.094 0.074 0.060 0.093 0.077 0.077
GeoT 0.042 0.046 0.059 0.055 0.046 0.050 0.073 0.053 0.053 0.062 0.070 0.071 0.080 0.075 0.049 0.107 0.083 0.074
Ours 0.043 0.037 0.061 0.051 0.044 0.029 0.048 0.044 0.045 0.061 0.068 0.068 0.088 0.071 0.054 0.091 0.091 0.074
A.4 Iterative Pose Refinement
We experiment with a simple iterative refinement scheme proposed in [44], where we can iteratively re-estimate the transformation by pruning out outliers.
$\mathbf{R,t}=\max_{\mathbf{R_{i}},\mathbf{t_{i}}}\sum_{i}{\llbracket\lVert\mathbf{R}\mathbf{x}_{i}+\mathbf{t}-\mathbf{y}_{i}\rVert^{2}<\tau_{a}\rrbracket},$ (12)
[…]
Refer to caption Refer to caption
Figure 7: Improvement in Reigstration Recall on increasing the number of iterations for pose refinement. We see that the performance saturates after the 5th iteration for 3DLoMatch and after 2 iterations for 3DMatch.
A.5 Qualitative Results
We provide more visual results in Fig. 8 and Fig. 9.
Figure 8: Qualitative Results on the ModelLoNet dataset. Row (a)-(d) show the cases with successful registration. While the rows (e) and (f) show the failure cases. The source cloud is shown in Green with the red color representing for the matched superpoints. Similarily target cloud is in red color with matching superpoints in green color. We see that our approach only fails in cases which are very ambigous.
Figure 9: Qualitative Results on the 3DLoMatch dataset. Row (a)-(d) show the cases with successful registration. While the rows (e) and (f) show the failure cases. The source cloud is shown in Green with the red color representing for the matched superpoints. Similarily target cloud is in red color with matching superpoints in green color.

### v4

@@ III-C Loss Functions
Transformation Loss. We apply the L1 loss on the predicted transformed locations of all keypoints with the predicted and ground truth transformation matrix. This is different from [59]. Since we do not predict corresponding points for each point cloud, we do not need to compute two transformation matrices and thus we do not have to force the network to learn that both matrices are inverses of each other.
@@
Overlap Loss. Inspired by [59], we estimate the overlap values $\mathbf{\hat{O}_{\mathbf{X^{\prime}}}}$ and $\mathbf{\hat{O}_{\mathbf{Y^{\prime}}}}$ using a separate MLP layer based on the enhanced feature $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ and $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$, respectively. The overlap estimation is formulated as a binary classification problem, so we use the binary cross-entropy loss:
$\mathcal{L}^{X}_{o}=-\frac{1}{M^{\prime}}\sum_{i}^{M^{\prime}}{o_{x,i}^{*}\cdot\log\hat{o}_{x,i}+(1-o_{x,i}^{*})\cdot\log\left(1-\hat{o}_{x,i}\right)},$ (4)
where $\hat{o}_{x,i}$ is the estimated overlap probability and $o_{x,i}^{*}$ is the ground truth probability. We compute the overlap loss $\mathcal{L}^{Y}_{o}$ for the target point cloud similarly.
@@
The final loss is a weighted sum of all the losses with
@@
IV Experiments
We evaluate our approach on three datasets with overlap ranging from 10% to 75%. The first dataset is on synthetic ModelNet dataset with two benchmarks settings following [21, 59]. The second one is 3DMatch [63] with two benchmarks following [21, 59, 39, 61]. The last one is on the challenging large-scale outdoor KITTI dataset [16]
IV-A Implementation details
Our approach is implemented using the PyTorch framework [36] on a system with an Intel i9-1300K CPU and a single RTX 3090 GPU. The network training is performed with the AdamW optimizer [31], using a learning rate of 0.0001 and a weight decay of 0.0001. We tain the network for 400 epochs with batch size of 4 on ModelNet, 50 epochs with batch size of 4 on 3DMatch, and 80 epochs with batch size of 1 on KITTI.
IV-B ModelNet and ModelLoNet Benchmarks
The ModelNet40 [54] dataset comprises of synthetic CAD models. Following the data setting in [21, 59], the point clouds are randomly sampled from mesh faces of the CAD models, cropped and subsampled.
Our network is trained exclusively on ModelNet, and evaluated for generalization on ModelLoNet. For benchmarking the performance of our model we use the Relative Rotation Error (RRE) and Relative translation Error (RTE) and Chamfer Distance (CD) as the primary metrics, following [59].
TABLE I: Registration results on ModelNet and ModelLoNet.
@@ IV-B ModelNet and ModelLoNet Benchmarks
PNLK [2] 29.725 0.297 0.0235 48.567 0.507 0.0367
OMNet [55] 2.947 0.032 0.0015 6.517 0.129 0.0074
DCPv2 [31] 11.975 0.171 0.0117 16.501 0.300 0.0268
RPMNet [58] 1.712 0.018 0.00085 7.342 0.124 0.0050
Predator [21] 1.739 0.019 0.00089 5.235 0.132 0.0083
RegTR [59] 1.473 0.014 0.00078 3.930 0.087 0.0037
GeoT [39] 1.568 0.018 - 3.809 0.102 -
@@
The results are shown in Table I. We compare against correspondence-based approaches [21, 59], coarse-to-fine registration approaches [39], and end-to-end methods [2, 58, 55]. Our approach performs well on both benchmarks improving results on ModelNet benchmark by 15% in RRE and 21% in RTE. The low chamfer error suggests that predicted correspondences have very high accuracy. Our approach is also able to outperform methods using post-processing steps like RANSAC [21] by a significant margin.
IV-C 3DMatch and 3DLoMatch Benchmarks
@@
TABLE II: Registration results on 3DMatch and 3DLoMatch.
@@
We compare our approach against several learned correspondence-based algorithms [18, 11, 4, 21] and coarse-to-fine approaches [61, 39]. The quantitative results are shown in Table II. For the 3DMatch benchmark, our approach outperforms all the previous methods in all but the FMR metric. This implies that in cases of significant overlap (>30%), the superpoint correspondences are very distinctive and accurate. For the 3DLoMatch benchmark, in comparison to all the approaches that do not use any post-processing, e.g. Predator[21] and CoFiNet[61], our method performs significantly better. GeoT [39] uses Local to Global Refinement (LGR) as the refinement step to get best results. In our case, outliers are filtered by the correlation weights, thus providing accurate correspondences. The validity of the correspondences obtained by our approach can be verified by comparing the mean Inlier Ratio (IR) in Table II. We get the highest mean IR, almost 19% better than the second-best approach on 3DMatch and 14% higher in 3DLoMatch. Our approach does not perform the best in FMR because we prioritize outputting correct correspondences over a large number of, yet potentially noisy matchings. In contrast, CofiNet [61], has a very high FMR but low IR due to its propensity for producing many inaccurate matches.
One of the problems associated with superpoint matching in point clouds is the resolution issue, where we might not have one-to-one correspondences due to subsampling. In our case, this is automatically handled by the correlation weights. The correlation weights are lower for weakly matching superpoints and thus even if there is no one-to-one correspondence, weak correspondences will be used according to their weights to compute the correct SE(3) transformation. Figure 3 shows the correlation weight distribution on the 3DMatch and 3DLoMatch benchmark. In 3DMatch, we have many high correlation weights, suggesting that many sample points are good correspondences. On 3DLoMatch, we have much fewer good correlations suggesting that many of the sampled superpoints are not good matches, which is expected because of very low overlap.
Refer to caption Refer to caption
Figure 3: Correlation weight distribution on 3DMatch. Here, many points exhibit moderate correlations, suggesting that a considerable portion of sample points are reasonably good matches. In contrast, 3DLoMatch has fewer points that exhibit strong correlations due to the low data overlap.
Figure 4: Qualitative Results on 3DMatch. The source cloud is shown in blue with the red color representing the matched superpoints. Similarily target cloud is shown in yellow color with matching superpoints in blue. (Best viewed in color.)
IV-D KITTI Benchmark
KITTI odometry consists of 11 sequences of outdoor driving scenarios scanned using the HDL64 LiDAR sensor. Following  [21], we use sequence 0-5 for training, 6-7 for validation, and 8-10 for testing. As in  [21, 39] we only use the point cloud pairs that are at least 10m apart and their ground-truth poses are refined using ICP for all evaluations. Table III shows the results obtained on KITTI. We can see that the baseline [59] performs significantly worse in RTE. Since the correspondences are predicted, they do not align one-on-one with actual correspondences, thus significantly reducing registration accuracy. We achieve state of the art results in RRE and RR while performing comparably in RTE. This showcases the effectiveness of our approach in large scale outdoor point clouds. Figure 1 shows an example of matched superpoints on KITTI dataset, our approach is able to find very distinctive correspondences to register the point cloud pair.
TABLE III: Registration results on KITTI.
[…]
3DFeat [57] 25.9 0.25 96.0
FCGF [42] 9.5 0.30 96.6
D3Feat [4] 7.2 0.30 99.8
Predator [21] 6.8 0.27 99.8
CofiNet [61] 8.2 0.41 99.8
RegTR [59] 49.9 0.61 99.1
GeoTransformer [39] 6.8 0.24 99.8
Unified BEV [27] 7.5 0.26 99.8
Ours 7.2 0.24 99.8
IV-E Ablation Studies
TABLE IV: Comparisons of different outlier filtering methods.
Outlier Filtering 3DMatch 3DLoMatch
Method RRE RTE RR RRE RTE RR Time
No filtering 1.462 0.046 93.4 2.652 0.074 64.6 0.073
Ours + RANSAC 2.788 0.090 82.3 4.653 0.115 32.0 0.141
Ours (top 15%) + RANSAC 1.701 0.051 93.0 2.696 0.078 64.5 0.153
@@ IV-E Ablation Studies
TABLE V: Effectiveness of different loss terms.
@@
Using correlation scores for filtering outliers. We study the effectiveness of outlier filtering in Table IV. In the ‘No filtering’ approach, we use all the superpoints with their scores without any filtering. A little surprisingly, it achieves very competitive results, likely because we use the correlation weights of the matchings in the transformation matrix estimation in Eq.(2). If the outlier filtering is used (the last row), the best accuracy can be obtained.
We also experiment with RANSAC on the set of superpoints correspondences with and without outlier filtering, respectively. Without any filtering (‘Ours + RANSAC’), the large number of outliers impose significant challenges for RANSAC to find the optimal solution. Even with filtering (‘Ours (top 15%) + RANSAC’), we can see that RANSAC gives slightly worse results, partially because it does not always converge to the optimal solution. Furthermore, RANSAC requires twice as much time for the estimation of the transformation matrix compared to our method.
@@
Effectiveness of the loss terms. We also analyze the effectiveness of each loss function. Table V shows the results with different loss function configurations. We see that by just using the feature loss along with the transformation loss, the model is able to achieve good performance. Using the overlap loss further helps the model prune remaining outliers.
Matching strategy. The choice of a matching strategy is crucial for precise pose estimation. In prior approaches, such as [39, 58, 61], the Sinkhorn algorithm has commonly been employed to compute optimal transport and derive matching scores for correspondences. In our research, we opt for Global Softmax, which, as evidenced by our experiments, outperforms the Sinkhorn approach (refer to Table VI). Note that while Sinkhorn algorithm requires careful parameter tuning, our approach is parameter-free and works out of the box. While Global Softmax has been previously used as a matching strategy in [49, 58], prior works only utilized it to identify matching points with probabilities exceeding a certain threshold, resulting in suboptimal outcomes. Our innovation lies in utilizing these probabilities as correspondence weights, facilitating gradient flow through the Weighted Kabsch-Umeyama Solver which leads to faster optimization and better results. We also experimented with Dual Softmax matching used in LofTR [46], but the difference in performance is less than 1% with Global Softmax being slightly better and faster than Dual-Softmax.
TABLE VI: Comparison of Matching Strategy
Methods 3DMatch 3DLoMatch Time
@@
Sinkhorn 1.509 0.046 89.2 2.622 0.074 59.6 0.094
Dual Softmax 1.441 0.045 93.6 2.589 0.074 65.1 0.080
MLP [59] 1.567 0.049 92.0 2.827 0.077 64.8 0.078
Global Softmax (Ours) 1.436 0.045 93.7 2.553 0.074 65.0 0.073
Data Augmentation Schemes Previous works have tried different data augmentation schemes.  [59] uses a weak data augmentation, perturbing poses by a small amount during training. While  [39, 58, 21] use large data augmentation, perturbing poses by a full range of motion. Weak data augmentation can slightly improve performance but deteriorates network generalization. Table VII shows that our approach achieves slightly better performance with weak data augmentation. But with strong data augmentation, the generalization on 3DLoMatch benchmarks improves RR by 1.8%.
TABLE VII: Comparison of Data Augmentation Schemes
Methods 3DMatch 3DLoMatch
@@
Ours (Weak Augmentation) 1.436 0.045 93.7 2.553 0.074 65.0
Ours (Large Augmentation) 1.472 0.045 93.4 2.611 0.074 66.8
V Conclusion
In this paper, we presented a strong baseline approach for point cloud registration. By using Global Softmax to directly match superpoint features, we remove the ambiguity of predicting non-existent corresponding points while using the softmax probabilities as correspondence weights allows us to filter outliers without any post-processing refinement. Experimental results on standard benchmarks show that our model achieves comparable or even better accuracy than state-of-the-art methods. We do not advocate our approach to be the solution for point cloud registration. Rather, we’d like to emphasize the role of matching strategy for point cloud registration, showcasing the feasibility of achieving high accuracy without cumbersome ad hoc postprocessing.